OpenStudy (anonymous):

The mean average temperature for Washington, DC is 53.8 degrees. The monthly average temperature ranges up to 32.5 degrees above and below this value. If t=0, represents March and the phase shift is -8. Write a model for the average monthly temperature in September. According to the model, what is the average temperature in September.

OpenStudy (anonymous):

A collection of quarters and nickels is worth $14.80. There are 88 coins in all. how many of each coin are there? quarters?? nickels??

OpenStudy (perl):

you can use y = A*cos[2Pi/Period*(t - p.s.)] + v.s. or y = A*sin[2Pi/Period*(t - p.s.)] + v.s.

OpenStudy (perl):

does the problem say to use sine or cosine ?

OpenStudy (perl):

I think we want to use cosine , since the phase shift -8 makes sense with cosine (using the fact that January is the coldest month in the northern hemisphere). y = 32.5 cos [ 2Pi/12 (t - (-8) ] + 53.8 simplified this gives y = 32.5 cos ( pi/6 ( t + 8) ) + 53.8 now can you answer the question.

OpenStudy (perl):

also be careful about what month september is

OpenStudy (anonymous):

I agree, cosine is the function to use.... However...I thought the model could be: y=32.5cos ( Pi/6t+8 ) +53.8... this differs from your model within the quantity...cos (Pi/6t+Pi/6(8) ).... What do you think ?? If t=0, represents March.... then September would be t=6 ????

OpenStudy (anonymous):

@Austin122131 Let x and y be the number of nickels and quarters respectively. {.05 x + .25 y = 14.80, x + y = 88} {x = 36., y = 52.}